Find the energy released in the alpha decay
238
92 U →234
90 T h +
4
2 He.
1
Expert's answer
2020-12-29T15:32:11-0500

mi = 238.050788 u

mf = 234.043601 u + 4.002603 u = 238.046204 u

Δm = |mf-mi| = "4.584 \times 10^{-3} \;u"

"E = Δmc^2 \\

E = 4.584 \times 10^{-3} \times 931.5 = 4.26 \;MeV"

Answer: 4.26 MeV